OpenStudy (anonymous):

1. Is (1,5) a solution to the equation y=x-2?

OpenStudy (aravindg):

Just plug in the numbers and see

OpenStudy (anonymous):

Does 1-2=5?

OpenStudy (aravindg):

y=x-2 put y=5 x=1 If equation is satisfied then it is a solution.

OpenStudy (anonymous):

so the answer is No

OpenStudy (kewlgeek555):

Yeah, like @AravindG said, plug in the numbers, but remember that 1 = x and y = 5 - don't get confused.

OpenStudy (anonymous):

So the answer is no

OpenStudy (anonymous):

the answer is no

OpenStudy (anonymous):

1-2 = -1

OpenStudy (anonymous):

Is (-2,-6) a solution to the equation y=3x

OpenStudy (anonymous):

plug it in yourself you can do it then ill tell you if you are right or wrong

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are we multiplying?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so yes

OpenStudy (anonymous):

the answer is yes it would look like this -6=3(-2)

OpenStudy (anonymous):

Ok hold on

OpenStudy (anonymous):

brb 20 mins